Two busea A and B leave the same bus depot, A towards the north and B towards the East. After four hours the distance between the two buses is 100 km. Determine the speed of the bus A if it travels at a speed of 5 km/hr more than that of bus B.
Answer by rothauserc(4718)

Use Pythagorean Theorem and rate * time = distance
:
(4x+20)^2 + (4x)^2 = 100^2
:
16x^2 +160x +400 +16x^2 = 10000
:
32x^2 +160x -9600 = 0
:
x^2 +5x -300 = 0
:
(x+20)*(x-15) = 0
:
we have two solutions for x, -20 and 15
we reject the negative solution
:
x = 15
:
A = 15 +5 = 20 km/hr
: